# Nonreciprocal heat transport in the Kitaev chiral spin liquid

### v1

@@ IV Discussion
The nonlinear conductivity is evaluated to be smaller than $\tilde{\kappa}^{(1)}$ by approximately the order of relaxation time ($\tilde{\kappa}^{(2)}/\tilde{\kappa}^{(1)}\sim\tau\times 10^{2}$ m/K). According to $\partial T/\partial X=160$ K/m [16, 28], each linear and nonlinear part of the thermal current (the first and second terms in Eq. (11)) is quantitatively estimated as $\tilde{\mathcal{J}}^{(1)}\sim 10^{4}~{}\mathrm{W/m^{2}}$ and $\tilde{\mathcal{J}}^{(2)}\sim 1~{}\mathrm{W/m^{2}}$. The nonreciprocal current is orders of magnitude smaller than the conventional contribution. However, there are several possible suggestion to enhance the nonlinear effect: (1) shorten the length of systems in the $Y$ direction, (2) increase the steepness of the thermal gradient, (3) use a sample with long relaxation time, and so on. The nonlinear effects, which are derived from the edge states, may be observed by focusing on the $Y$-directional length dependence of the samples. Additionally, measuring the dependence of $\tilde{\mathcal{J}}/(-\frac{\partial T}{\partial X})$ on the thermal gradient would enable us to estimate the conductivities on the basis of the intercept and slope.
@@ V Conclusion
We have theoretically examined nonreciprocal heat transport in the Kitaev chiral spin liquid considering open edges. The DMI at the edge induces an asymmetry in the Majorana band, thereby leading to nonreciprocal heat transport. We stress that the nonreciprocal transport considered here occurs even if we take into account the contributions from both upper and lower edges. This feature is contrasted to the previous studies which focus on one particular edge [8]. This nonreciprocal heat transport is enhanced in the low temperature region, while the dilute flux excitation in the bulk can further enhance its signature. Furthermore, there exists the potential to manipulate both of direction and magnitude of the nonreciprocal heat current by varying the external magnetic fields, corresponding to the sign and the absolute value of $D_{1}=D_{3}=h_{x}D_{x}-h_{y}D_{y}$. Further development of nonreciprocal phenomena is expected in various Majorana systems.
Figure 8: All perturbation processes of DMI at the upper edge. The red and blue bonds represents flips of the $Z_{2}$ gauge fields by DMI and by Zeeman term.
Figure 9: All perturbation processes of DMI at the lower edge. The red and blue bonds represents flips of the $Z_{2}$ gauge fields by DMI and by Zeeman term.
@@ Acknowledgements
The authors are grateful to T. Morimoto for invaluable discussion. M. O. T. is supported by a Japan Society for the Promotion of Science (JSPS) Fellowship for Young Scientists and by Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program.” M. G. Y. is supported by a Japan Science and Technology Agency (JST) PRESTO (Grant No. JPMJPR225B) and by the Center of Innovation for Sustainable Quantum AI (JST Grant No. JPMJPF2221). This work is supported by JST CREST Grant No. JPMJCR19T5 and JSPS KAKENHI No. JP22H01221, No. JP22J20066, No. JP22K14005 and No. JP23K20828.
Appendix A Perturbation of DMI

### v2

@@ IV Discussion
The nonlinear conductivity is evaluated to be smaller than $\tilde{\kappa}^{(1)}$ by approximately the order of relaxation time ($\tilde{\kappa}^{(2)}/\tilde{\kappa}^{(1)}\sim\tau\times 10^{2}$ m/K). According to $\partial T/\partial X=160$ K/m [19, 37], the ratio of the conductivities is approximated as $\tilde{\kappa}^{(2)}\left(-\frac{\partial T}{\partial X}\right)/\tilde{\kappa}^{(1)}\sim 10^{-4}$, and each linear and nonlinear component of the thermal current (the first and second terms in Eq. (11)) is quantitatively estimated as $\tilde{\mathcal{J}}^{(1)}\sim 10^{4}~{}\mathrm{W/m^{2}}$ and $\tilde{\mathcal{J}}^{(2)}\sim 1~{}\mathrm{W/m^{2}}$. Although the nonreciprocal current is orders of magnitude smaller than the conventional contribution, this nonreciprocal characteristic would be observable, as currently accessible experimental techniques enable us to detect $\kappa_{xx}\sim 1~{}\mathrm{W/Km}$ and $\kappa_{xy}\sim 10^{-4}~{}\mathrm{W/Km}$ for the Kitaev candidate $\mathrm{\alpha}\text{-}\mathrm{RuCl_{3}}$ at $5~{}\mathrm{K}$ [19, 20]. Also, there are several possible suggestion to enhance the nonlinear effect: (1) shorten the length of systems in the $Y$ direction, (2) increase the steepness of the thermal gradient, (3) use a sample with long relaxation time, and so on. The nonlinear effects, which are derived from the edge states, may be observed by focusing on the $Y$-directional length dependence of the samples. Additionally, measuring the dependence of $\tilde{\mathcal{J}}/(-\frac{\partial T}{\partial X})$ on the thermal gradient would enable us to estimate the conductivities on the basis of the intercept and slope.
Figure 8: All perturbation processes of the DMI at the upper edge. The red and blue bonds represents flips of the $Z_{2}$ gauge fields by the DMI and by Zeeman term.
Figure 9: All perturbation processes of the DMI at the lower edge. The red and blue bonds represents flips of the $Z_{2}$ gauge fields by the DMI and by Zeeman term.
@@ V Conclusion
We have theoretically examined nonreciprocal heat transport in the Kitaev chiral spin liquid considering open edges. The DMI at the edge induces an asymmetry in the Majorana band, thereby leading to nonreciprocal heat transport. We stress that the nonreciprocal transport considered here occurs even if we take into account the contributions from both upper and lower edges. This feature is contrasted to the previous studies which focus on one particular edge [11]. This nonreciprocal heat transport is enhanced in the low temperature region, while the dilute flux excitation in the bulk can further enhance its signature. Furthermore, there exists the potential to manipulate both of direction and magnitude of the nonreciprocal heat current by varying the external magnetic fields, corresponding to the sign and the absolute value of $D_{1}=D_{3}=h_{x}D_{x}-h_{y}D_{y}$. As a future perspective, it would be interesting to explore how an inhomogeneous potential influences the bulk thermal current [38]. Further development of nonreciprocal phenomena is expected in various Majorana systems.
@@ Acknowledgements
The authors are grateful to T. Morimoto and Y. Kasahara for invaluable discussion. M. O. T. is supported by a Japan Society for the Promotion of Science (JSPS) Fellowship for Young Scientists and by Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program.” M. G. Y. is supported by a Japan Science and Technology Agency (JST) PRESTO (Grant No. JPMJPR225B) and by the Center of Innovation for Sustainable Quantum AI (JST Grant No. JPMJPF2221). This work is supported by JST CREST Grant No. JPMJCR19T5 and JSPS KAKENHI No. 23K20828, No. JP22H01221, No. JP22J20066 and No. JP22K14005.
Appendix A Perturbation of the DMI
@@
Figure 10: The band structures when (a) $\bm{h}=(0.1,0.1,0.1)$, $\bm{D}=(0,0,0)$; (b) $\bm{h}=(0.1,-0.1,0.1)$, $\bm{D}=(0,0,0)$; (c) $\bm{h}=(0.2,-0.2,0.2)$, $\bm{D}=(0.1,0.1,0.1)$; and (d)$\bm{h}=(-0.2,0.2,0.2)$, $\bm{D}=(0.1,0.1,0.1)$. (a) and (b), cases without the DMI, are numerically identical and symmetric. (c) and (d) are asymmetric and inverted.
Figure 11: In a variety of parameters, (a) the linear conductivity $\kappa^{(1)}$, (b) the nonlinear conductivity $\kappa^{(2)}$. Each of blue, red, green and purple lines represents $\bm{h}=(0.1,0.1,0.1)$, $\bm{D}=(0,0,0)$; $\bm{h}=(0.1,-0.1,0.1)$, $\bm{D}=(0,0,0)$; $\bm{h}=(0.2,-0.2,0.2)$, $\bm{D}=(0.1,0.1,0.1)$; and $\bm{h}=(-0.2,0.2,0.2)$, $\bm{D}=(0.1,0.1,0.1)$. The inset shows the nonlinear conductivity at the extreme low temperature.
Appendix B Band structures and heat conductivities on various parameters
In this section, we present band structures and conductivities for various sets of a magnetic field $\bm{h}$ and the DMI $\bm{D}$, particularly for the cases where $\bm{h}=(0.1,0.1,0.1)$, $\bm{D}=(0,0,0)$; $\bm{h}=(0.1,-0.1,0.1)$, $\bm{D}=(0,0,0)$; $\bm{h}=(0.2,-0.2,0.2)$, $\bm{D}=(0.1,0.1,0.1)$; and $\bm{h}=(-0.2,0.2,0.2)$, $\bm{D}=(0.1,0.1,0.1)$.
In Figs. 10(a) and 10(b), we show the band structures with different magnetic fields but without the DMI, which are numerically identical and symmetric. In Figs. 10(c) and 10(d), we show the effects of stronger magnetic fields, compared with Figs. 3(a) and 3(b), on band structures. The increased magnetic field intensifies the asymmetry and widens the gap at $k_{X}=\pm\pi$. Notably, the transformation “$h_{x}\rightarrow-h_{x}$ and $h_{y}\rightarrow-h_{y}$” inverts the band structures, here again.
In Figs. 11(a) and 11(b), we show the linear and nonlinear conductivities. For the blue and red line, i.e., in the case of $\bm{D}=0$, the liner conductivities retain as same magnitude as those in Fig. 5(a), and the nonlinear conductivities exhibit no nonzero value. In contrast, for the green and purple lines, both the linear and nonlinear conductivities are enhanced by factors on the order of $10$ to $10^{2}$, compared with those in Figs. 5(a) and 5(b). Furthermore, for the large magnetic field, the temperature dependence of the heat conductivities at low temperatures are qualitatively different from the case with smaller magnetic field.